OpenStudy (anonymous):

you are riding your bicycle to prepare for a race.It takes you 12min to go 2.5 miles if you continue traveling at the same rate how long will it take you to got 7 miles

OpenStudy (ranga):

2.5 miles takes 12 minutes. 1 mile will take 12/2.5 minutes 7 miles will take 7 * 12/2.5 minutes = 33.6 minutes.

OpenStudy (ranga):

Another way to solve this problem is by direct proportions. 12min to go 2.5 miles. The rate is 2.5 / 12 miles per minute. Assume it takes x minutes to to travel 7 miles. The rate is 7 / x miles per minute. The person rides the bicycle at the same rate. Therefore, 2.5/12 = 7 / x cross multiply: 2.5x = 7*12 divide both sides by 2.5 x = 7*12/2.5 = 33.6 minutes.